# On the Identification of Temporally Causal Representation with Instantaneous Dependence

### v2

@@ 6 Conclusion
Let $\hat{c}_{t,k},\hat{c}_{t,l}$ be two different variables that are not adjacent in the estimated Markov network $\mathcal{M}_{\hat{\mathbf{c}}_{t}}$ over $\hat{{\mathbf{c}}}_{t}=\{\hat{{\mathbf{z}}}_{t-1},\hat{{\mathbf{z}}}_{t}\}$. We conduct the first-order derivative w.r.t. $\hat{c}_{t,k}$ and have
$\frac{\partial\log p(\hat{{\mathbf{c}}}_{t}|\hat{\mathbf{z}}_{t-2})}{\partial\hat{c}_{t,k}}=\sum_{i=1}^{2n}\frac{\partial\log p({{\mathbf{c}}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,i}}\cdot\frac{\partial c_{t,i}}{\partial\hat{c}_{t,k}}+\frac{\partial\log|\mathbf{J}_{h_{c},t}|}{\partial\hat{c}_{t,k}}.$ (19)
We further conduct the second-order derivative w.r.t. $\hat{c}_{t,k}$ and $\hat{c}_{t,l}$, then we have:
$\begin{split}\frac{\partial\log p(\hat{{\mathbf{c}}}_{t}|\hat{\mathbf{z}}_{t-2})}{\partial\hat{c}_{t,k}\partial\hat{c}_{t,l}}&=\sum_{i=1}^{2n}\sum_{j=1}^{2n}\frac{\partial^{2}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,i}\partial c_{t,j}}\cdot\frac{\partial c_{t,i}}{\partial\hat{c}_{t,k}}\cdot\frac{\partial c_{t,j}}{\partial\hat{c}_{t,l}}\\ &+\sum_{i=1}^{2n}\frac{\partial\log p({{\mathbf{c}}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,i}}\cdot\frac{\partial^{2}c_{t,i}}{\partial\hat{c}_{t,k}\partial\hat{c}_{t,l}}+\frac{\partial\log|\mathbf{J}_{h_{c},t}|}{\partial\hat{c}_{t,k}\partial\hat{c}_{t,l}}.\end{split}$ (20)
Since $\hat{c}_{t,k},\hat{c}_{t,l}$ are not adjacent in $\mathcal{M}_{\hat{\mathbf{c}}_{t}}$, $\hat{c}_{t,k}$ and $\hat{c}_{t,l}$ are conditionally independent given $\hat{{\mathbf{c}}}_{t}\textbackslash\{\hat{c}_{t,k},\hat{c}_{t,l}\}$. Utilizing the fact that conditional independence can lead to zero cross derivative [38], for each value of $\hat{{\mathbf{z}}}_{t-2}$, we have
$\begin{split}\frac{\partial^{2}\log p(\hat{{\mathbf{c}}}_{t}|\hat{{\mathbf{z}}}_{t-2})}{\partial\hat{c}_{t,k}\partial\hat{c}_{t,l}}=&\frac{\partial^{2}\log p(\hat{c}_{t,k}|\hat{{\mathbf{c}}}_{t}\textbackslash\{\hat{c}_{t,k},\hat{c}_{t,l}\},\hat{{\mathbf{z}}}_{t-2})}{\partial\hat{c}_{t,k}\partial\hat{c}_{t,l}}+\frac{\partial^{2}\log p(\hat{c}_{t,l}|{\mathbf{c}}_{t}\textbackslash\{\hat{c}_{t,k},\hat{c}_{t,l}\},\hat{{\mathbf{z}}}_{t-2})}{\partial\hat{c}_{t,k}\partial\hat{c}_{t,l}}\\ &+\frac{\partial^{2}\log p(\hat{\mathbf{c}}_{t}\textbackslash\{\hat{c}_{t,k},\hat{c}_{t,l}\}|\hat{{\mathbf{z}}}_{t-2})}{\partial\hat{c}_{t,k}\partial\hat{c}_{t,l}}=0.\end{split}$ (21)
Bring in Equation 21, Equation 20 can be further derived as
$\begin{split}0&=\underbrace{\sum_{i=1}^{2n}\frac{\partial^{2}\log p({{\mathbf{c}}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,i}^{2}}\cdot\frac{\partial c_{t,i}}{\partial\hat{c}_{t,k}}\cdot\frac{\partial c_{t,i}}{\partial\hat{c}_{t,l}}}_{\textbf{(i) }i=j}+\underbrace{\sum_{i=1}^{2n}\sum_{(j,i)\in\mathcal{E}(\mathcal{M}_{{\mathbf{c}}_{t}})}\frac{\partial^{2}\log p({{\mathbf{c}}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,i}\partial c_{t,j}}\cdot\frac{\partial c_{t,i}}{\partial\hat{c}_{t,k}}\cdot\frac{\partial c_{t,j}}{\partial\hat{c}_{t,l}}}_{\textbf{(ii)}c_{t,i}\text{ and }c_{t,j}\text{ are adjacent in $\mathcal{M}_{{\mathbf{c}}_{t}}$}}\\ &+\underbrace{\sum_{i=1}^{2n}\sum_{(j,i)\notin\mathcal{E}(\mathcal{M}_{{\mathbf{c}}_{t}})}\frac{\partial^{2}\log p({{\mathbf{c}}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,i}\partial c_{t,j}}\cdot\frac{\partial c_{t,i}}{\partial\hat{c}_{t,k}}\cdot\frac{\partial c_{t,j}}{\partial\hat{c}_{t,l}}}_{\textbf{(iii)}c_{t,i}\text{ and }c_{t,j}\text{ are {not} adjacent in $\mathcal{M}_{{\mathbf{c}}_{t}}$}}\\ &+\sum_{i=1}^{2n}\frac{\partial\log p({{\mathbf{c}}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,i}}\cdot\frac{\partial^{2}c_{t,i}}{\partial\hat{c}_{t,k}\partial\hat{c}_{t,l}}+\frac{\partial\log|\mathbf{J}_{h_{c},t}|}{\partial\hat{c}_{t,k}\partial\hat{c}_{t,l}}.\end{split}$ (22)
@@
Similar to Equation 21, we have $\frac{\partial^{2}p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,i}\partial c_{t,j}}=0$ when $c_{t,i},c_{t,j}$ are not adjacent in $\mathcal{M}_{{\mathbf{c}}_{t}}$. Thus, Equation 22 can be rewritten as
$\begin{split}0=&\sum_{i=1}^{2n}\frac{\partial^{2}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,i}^{2}}\cdot\frac{\partial c_{t,i}}{\partial\hat{c}_{t,k}}\cdot\frac{\partial c_{t,i}}{\partial\hat{c}_{t,l}}+\sum_{i=1}^{2n}\sum_{(j,i)\in\mathcal{E}(\mathcal{M}_{{\mathbf{c}}})}\frac{\partial^{2}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,i}\partial c_{t,j}}\cdot\frac{\partial c_{t,i}}{\partial\hat{c}_{t,k}}\cdot\frac{\partial c_{t,j}}{\partial\hat{c}_{t,l}}\\ &+\sum_{i=1}^{2n}\frac{\partial\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,i}}\cdot\frac{\partial^{2}c_{t,i}}{\partial\hat{c}_{t,k}\partial\hat{c}_{t,l}}+\frac{\partial\log|\mathbf{J}_{h_{c},t}|}{\partial\hat{c}_{t,k}\partial\hat{c}_{t,l}}.\end{split}$ (23)
@@
Then for each $m=1,2,\cdots,n$ and each value of $z_{t-2,m}$, we conduct partial derivative on both sides of Equation 23 and have:
@@
$\begin{split}0=&\sum_{i=1}^{2n}\frac{\partial^{3}\log p({{\mathbf{c}}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,i}^{2}\partial z_{t-2,m}}\cdot\frac{\partial c_{t,i}}{\partial\hat{c}_{t,k}}\cdot\frac{\partial c_{t,i}}{\partial\hat{c}_{t,l}}+\sum_{i=1}^{2n}\sum_{(j,i)\in\mathcal{E}(\mathcal{M}_{{\mathbf{c}}})}\frac{\partial^{3}\log p({{\mathbf{c}}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,i}\partial c_{t,j}\partial z_{t-2,m}}\cdot\frac{\partial c_{t,i}}{\partial\hat{c}_{t,k}}\cdot\frac{\partial c_{t,j}}{\partial\hat{c}_{t,l}}\\ &+\sum_{i=1}^{2n}\frac{\partial^{2}\log p(c_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,i}\partial z_{t-2,m}}\cdot\frac{\partial c_{t,i}^{2}}{\partial\hat{c}_{t,k}\partial\hat{c}_{t,l}}\end{split},$ (24)
@@
According to Assumption A2, we can construct $4n+2|\mathcal{M}_{{\mathbf{c}}}|$ different equations with different values of $z_{t-2,m}$, and the coefficients of the equation system they form are linearly independent. To ensure that the right-hand side of the equations are always 0, the only solution is
$\begin{split}\frac{\partial c_{t,i}}{\partial\hat{c}_{t,k}}\cdot\frac{\partial c_{t,i}}{\partial\hat{c}_{t,l}}=&0,\end{split}$ (25)
$\begin{split}\frac{\partial c_{t,i}}{\partial\hat{c}_{t,k}}\cdot\frac{\partial c_{t,j}}{\partial\hat{c}_{t,l}}=&0,\end{split}$ (26)
$\begin{split}\frac{\partial c_{t,i}^{2}}{\partial\hat{c}_{t,k}\partial\hat{c}_{t,l}}=&0.\end{split}$ (27)
@@
According to the aforementioned results, for any two different entries $\hat{c}_{t,k},\hat{c}_{t,l}\in\hat{{\mathbf{c}}}_{t}$ that are not adjacent in the Markov network $\mathcal{M}_{\hat{{\mathbf{c}}}_{t}}$ over estimated $\hat{{\mathbf{c}}}_{t}$, we draw the following conclusions.
@@
(i) Equation 25 implies that, each ground-truth latent variable $c_{t,i}\in{\mathbf{c}}_{t}$ is a function of at most one of $\hat{c}_{t,k}$ and $\hat{c}_{t,l}$,
@@
(ii) Equation 26 implies that, for each pair of ground-truth latent variables $c_{t,i}$ and $c_{t,j}$ that are adjacent in $\mathcal{M}_{{\mathbf{c}}_{t}}$ over ${\mathbf{c}}_{t}$, they can not be a function of $\hat{c}_{t,k}$ and $\hat{c}_{t,l}$ respectively.
A2.2 Extension to Multiple Lags and Sequence Lengths
@@
For the sake of simplicity, we consider only one special case with $\tau=1$ and $L=2$ in Theorem A2. Our identifiability theorem can be actually extended to arbitrary lags and subsequences easily. For any given $\tau$, and subsequence which is centered at ${\mathbf{z}}_{t}$ with previous $lo$ and following $hi$ steps, i.e., ${\mathbf{c}}_{t}=\{{\mathbf{z}}_{t-lo},\cdots,{\mathbf{z}}_{t},\cdots,{\mathbf{z}}_{t+hi}\}$. In this case, the vector function $w(i,j,m)$ in Sufficient Variability Assumption should be modified as
@@
$\small\begin{split}w(i,j,m)=&\Big{(}\frac{\partial^{3}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-lo-1},\cdots,{\mathbf{z}}_{t-lo-\tau})}{\partial c_{t,1}^{2}\partial z_{t-lo-1,m}},\cdots,\frac{\partial^{3}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-lo-1},\cdots,{\mathbf{z}}_{t-lo-\tau})}{\partial c_{t,2n}^{2}\partial z_{t-lo-1,m}}\Big{)}\oplus\\ &\Big{(}\frac{\partial^{2}\log p(c_{t}|{\mathbf{z}}_{t-lo-1},\cdots,{\mathbf{z}}_{t-lo-\tau})}{\partial c_{t,1}\partial z_{t-lo-1,m}},\cdots,\frac{\partial^{2}\log p(c_{t}|{\mathbf{z}}_{t-lo-1},\cdots,{\mathbf{z}}_{t-lo-\tau})}{\partial c_{t,2n}\partial z_{t-lo-1,m}}\Big{)}\oplus\\ &\Big{(}\frac{\partial^{3}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-lo-1},\cdots,{\mathbf{z}}_{t-lo-\tau})}{\partial c_{t,i}\partial c_{t,j}\partial z_{t-lo-1,m}}\Big{)}_{(i,j)\in\mathcal{E}(\mathcal{M}_{{\mathbf{c}}_{t}})}.\end{split}$ (28)
@@
Besides, $2*(lo+hi+1)+2|\mathcal{M}_{{\mathbf{c}}_{t}}|$ values of linearly independent vector functions in $z_{t^{\prime},m}$ for $t^{\prime}\in[t-lo-1,\cdots,t-lo-\tau]$ and $m\in[1,\cdots,n]$ are required as well. The rest part of the theorem remains the same, and the proof can be easily extended in such a setting.
A2.3 Identifiablity of Latent Variables

### v3

@@ B.1 Relationships between Ground-truth and Estimated Latent Variables
Let $\hat{z}_{t,k},\hat{z}_{t,l}$ be two different variables that are not adjacent in the estimated Markov network $\mathcal{M}_{\hat{\mathbf{z}}_{t}}$ over $\hat{{\mathbf{z}}}_{t}=$. We conduct the first-order derivative w.r.t. $\hat{z}_{t,k}$ and have
$\frac{\partial\log p(\hat{{\mathbf{z}}}_{t}|\hat{\mathbf{z}}_{t-1})}{\partial\hat{z}_{t,k}}=\sum_{i=1}^{n}\frac{\partial\log p({{\mathbf{z}}}_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,i}}\cdot\frac{\partial z_{t,i}}{\partial\hat{z}_{t,k}}+\frac{\partial\log|\mathbf{J}_{h_{z},t}|}{\partial\hat{z}_{t,k}}.$ (A7)
We further conduct the second-order derivative w.r.t. $\hat{z}_{t,k}$ and $\hat{z}_{t,l}$, then we have:
$\begin{split}\frac{\partial^{2}\log p(\hat{{\mathbf{z}}}_{t}|\hat{\mathbf{z}}_{t-1})}{\partial\hat{z}_{t,k}\partial\hat{z}_{t,l}}&=\sum_{i=1}^{n}\sum_{j=1}^{n}\frac{\partial^{2}\log p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,i}\partial z_{t,j}}\cdot\frac{\partial z_{t,i}}{\partial\hat{z}_{t,k}}\cdot\frac{\partial z_{t,j}}{\partial\hat{z}_{t,l}}\\ &+\sum_{i=1}^{2}\frac{\partial\log p({{\mathbf{z}}}_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,i}}\cdot\frac{\partial^{2}z_{t,i}}{\partial\hat{z}_{t,k}\partial\hat{z}_{t,l}}+\frac{\partial^{2}\log|\mathbf{J}_{h_{z},t}|}{\partial\hat{z}_{t,k}\partial\hat{z}_{t,l}}.\end{split}$ (A8)
[…]
Bring in Equation (A9), Equation (A8) can be further derived as
[…]
@@
where $(j,i)\in\mathcal{E}(\mathcal{M}_{{\mathbf{z}}_{t}})$ denotes that $z_{t,i}$ and $z_{t,j}$ are adjacent in $\mathcal{M}_{{\mathbf{z}}_{t}}$. Similar to Equation (A9), we have $\frac{\partial^{2}p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,i}\partial z_{t,j}}=0$ when $z_{t,i},z_{t,j}$ are not adjacent in $\mathcal{M}_{{\mathbf{z}}_{t}}$. Thus, Equation (A10) can be rewritten as
$\begin{split}0=&\sum_{i=1}^{n}\frac{\partial^{2}\log p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,i}^{2}}\cdot\frac{\partial z_{t,i}}{\partial\hat{z}_{t,k}}\cdot\frac{\partial z_{t,i}}{\partial\hat{z}_{t,l}}+\sum_{i=1}^{n}\sum_{j:(j,i)\in\mathcal{E}(\mathcal{M}_{{\mathbf{z}}})}\frac{\partial^{2}\log p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,i}\partial z_{t,j}}\cdot\frac{\partial z_{t,i}}{\partial\hat{z}_{t,k}}\cdot\frac{\partial z_{t,j}}{\partial\hat{z}_{t,l}}\\ &+\sum_{i=1}^{n}\frac{\partial\log p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,i}}\cdot\frac{\partial^{2}z_{t,i}}{\partial\hat{z}_{t,k}\partial\hat{z}_{t,l}}+\frac{\partial\log|\mathbf{J}_{h_{z},t}|}{\partial\hat{z}_{t,k}\partial\hat{z}_{t,l}}.\end{split}$ (A11)
@@
Then for each $m=1,2,\cdots,n$ and each value of $z_{t-1,m}$, we conduct partial derivative on both sides of Equation (A11) and have:
$\begin{split}0=&\sum_{i=1}^{n}\frac{\partial^{3}\log p({{\mathbf{z}}}_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,i}^{2}\partial z_{t-1,m}}\cdot\frac{\partial z_{t,i}}{\partial\hat{z}_{t,k}}\cdot\frac{\partial z_{t,i}}{\partial\hat{z}_{t,l}}+\sum_{i=1}^{n}\sum_{j:(j,i)\in\mathcal{E}(\mathcal{M}_{{\mathbf{z}}})}\frac{\partial^{3}\log p({{\mathbf{z}}}_{t}|{\mathbf{z}}_{t-2})}{\partial z_{t,i}\partial z_{t,j}\partial z_{t-1,m}}\cdot\frac{\partial z_{t,i}}{\partial\hat{z}_{t,k}}\cdot\frac{\partial z_{t,j}}{\partial\hat{z}_{t,l}}\\ &+\sum_{i=1}^{n}\frac{\partial^{2}\log p(z_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,i}\partial z_{t-1,m}}\cdot\frac{\partial z_{t,i}^{2}}{\partial\hat{z}_{t,k}\partial\hat{z}_{t,l}}\end{split},$ (A12)
Finally we have
@@
$\begin{split}0=&\sum_{i=1}^{n}\frac{\partial^{3}\log p({{\mathbf{z}}}_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,i}^{2}\partial z_{t-1,m}}\cdot\frac{\partial z_{t,i}}{\partial\hat{z}_{t,k}}\cdot\frac{\partial z_{t,i}}{\partial\hat{z}_{t,l}}+\sum_{i=1}^{n}\frac{\partial^{2}\log p(z_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,i}\partial z_{t-1,m}}\cdot\frac{\partial z_{t,i}^{2}}{\partial\hat{z}_{t,k}\partial\hat{z}_{t,l}}\\ &+\sum_{i,j:(j,i)\in\mathcal{E}(\mathcal{M}_{{\mathbf{z}}})}\frac{\partial^{3}\log p({{\mathbf{z}}}_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,i}\partial z_{t,j}\partial z_{t-1,m}}\cdot\left(\frac{\partial z_{t,i}}{\partial\hat{z}_{t,k}}\cdot\frac{\partial z_{t,j}}{\partial\hat{z}_{t,l}}+\frac{\partial z_{t,j}}{\partial\hat{z}_{t,k}}\cdot\frac{\partial z_{t,i}}{\partial\hat{z}_{t,l}}\right)\end{split}.$ (A13)
@@
According to Assumption A2, we can construct $4n+|\mathcal{M}_{{\mathbf{c}}}|$ different equations with different values of $z_{t-1,m}$, and the coefficients of the equation system they form are linearly independent. To ensure that the right-hand side of the equations is always 0, the only solution is
$\begin{split}\frac{\partial z_{t,i}}{\partial\hat{z}_{t,k}}\cdot\frac{\partial z_{t,i}}{\partial\hat{z}_{t,l}}=&0,\end{split}$ (A14)
$\begin{split}\frac{\partial z_{t,i}}{\partial\hat{z}_{t,k}}\cdot\frac{\partial z_{t,j}}{\partial\hat{z}_{t,l}}+\frac{\partial z_{t,j}}{\partial\hat{z}_{t,k}}\cdot\frac{\partial z_{t,i}}{\partial\hat{z}_{t,l}}=&0,\end{split}$ (A15)
$\begin{split}\frac{\partial z_{t,i}^{2}}{\partial\hat{z}_{t,k}\partial\hat{z}_{t,l}}=&0.\end{split}$ (A16)
Bringing Eq A14 into Eq A15, at least one product must be zero, thus the other must be zero as well. That is,
$\begin{split}\frac{\partial z_{t,i}}{\partial\hat{z}_{t,k}}\cdot\frac{\partial z_{t,j}}{\partial\hat{z}_{t,l}}=&0.\end{split}$ (A17)
@@
According to the aforementioned results, for any two different entries $\hat{z}_{t,k},\hat{z}_{t,l}\in\hat{{\mathbf{z}}}_{t}$ that are not adjacent in the Markov network $\mathcal{M}_{\hat{{\mathbf{z}}}_{t}}$ over estimated $\hat{{\mathbf{z}}}_{t}$, we draw the following conclusions.
@@
(i) Equation (A14) implies that, each ground-truth latent variable $z_{t,i}\in{\mathbf{z}}_{t}$ is a function of at most one of $\hat{z}_{t,k}$ and $\hat{z}_{t,l}$,
@@
(ii) Equation (A17) implies that, for each pair of ground-truth latent variables $z_{t,i}$ and $z_{t,j}$ that are adjacent in $\mathcal{M}_{{\mathbf{z}}_{t}}$ over ${\mathbf{z}}_{t}$, they can not be a function of $\hat{z}_{t,k}$ and $\hat{z}_{t,l}$ respectively.
B.2 Extension to Multiple Lags and Sequence Lengths
@@
For the sake of simplicity, we consider only one special case with $\tau=1$ and $L=2$ in Theorem A2. Our identifiability theorem can be actually extended to arbitrary lags and subsequences easily. For any given $\tau$, and subsequence which is centered at ${\mathbf{z}}_{t}$ with previous $lo$ and following $hi$ steps, i.e., ${\mathbf{z}}_{t}=\{{\mathbf{z}}_{t-lo},\cdots,{\mathbf{z}}_{t},\cdots,{\mathbf{z}}_{t+hi-1}\}\in\mathbb{R}^{(lo+hi)\times n}$. In this case, the vector function $w(i,j,m)$ in Sufficient Variability Assumption should be modified as
@@
$\begin{split}w(i,j,m)=&\Big(\frac{\partial^{3}\log p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-lo-1},\cdots,{\mathbf{z}}_{t-lo-\tau})}{\partial z_{t,1}^{2}\partial z_{t-lo-1,m}},\cdots,\frac{\partial^{3}\log p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-lo-1},\cdots,{\mathbf{z}}_{t-lo-\tau})}{\partial z_{t,n}^{2}\partial z_{t-lo-1,m}}\Big)\oplus\\ &\Big(\frac{\partial^{2}\log p(z_{t}|{\mathbf{z}}_{t-lo-1},\cdots,{\mathbf{z}}_{t-lo-\tau})}{\partial z_{t,1}\partial z_{t-lo-1,m}},\cdots,\frac{\partial^{2}\log p(z_{t}|{\mathbf{z}}_{t-lo-1},\cdots,{\mathbf{z}}_{t-lo-\tau})}{\partial z_{t,2n}\partial z_{t-lo-1,m}}\Big)\oplus\\ &\Big(\frac{\partial^{3}\log p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-lo-1},\cdots,{\mathbf{z}}_{t-lo-\tau})}{\partial z_{t,i}\partial z_{t,j}\partial z_{t-lo-1,m}}\Big)_{(i,j)\in\mathcal{E}(\mathcal{M}_{{\mathbf{z}}_{t}})}.\end{split}$ (A18)
@@
Besides, $2\times n\times(lo+hi+1)+|\mathcal{M}_{{\mathbf{c}}_{t}}|$ values of linearly independent vector functions in $z_{t^{\prime},m}$ for $t^{\prime}\in[t-lo-1,\cdots,t-lo-\tau]$ and $m\in[1,\cdots,n]$ are required as well. The rest of the theorem remains the same, and the proof can be easily extended in such a setting.
B.3 Identifiability of Latent Variables